media%2Fa2a%2Fa2a496e8-ea71-4136-b8db-31


A project that provides annual cash flows of $15400 for nine years costs $67,000 today What is the NPV if the required return is 8 percent? What if it's 20 percent? At what discount rate would you be indifferent between accepting the project and rejecting it? 15,400 9 67,000 8% 20% Annual cash flows # of years Costs Required Return Required Return S 10 12 13 14 15 16 17 Complete the following analysis. Do not hard code values in your calculations. You must use the built-in Excel function to answer this question. NPV at 8 percent NPV at 20 percent IRR
Answer #1

In cell D15 enter the below formula
=PV(D9,D7,-D6,0)-D8

In cell D17 enter the below formula
=PV(D10,D7,-D6,0)-D8

In cell D19 enter the below formula
=RATE(D7,-D6,D8,0)